# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ 7.2 Validity
Validity, another cornerstone principle of psychometric validation, determines whether a test accurately measures its intended construct. Recent research has systematically investigated multiple aspects of validity in LLM psychometrics, addressing both methodological challenges and potential solutions unique to this emerging field.
@@ 7.2.1 Content Validity
Content validity ensures that a test comprehensively covers the construct it aims to measure. Major challenges of content validity include data contamination of well-established tests and novel items that are inadequately examined.
Data Contamination.
Some studies naively transfer human tests to LLMs. Given that LLMs are trained on vast Internet-scale corpora, it is highly probable that they have been exposed to instances of test items or similar content [Hagendorff et al., 2024, Jiang et al., 2024c]. This exposure can result in the reproduction of known token patterns during evaluation, which leads to inflated performance metrics or biases the reflection of the LLM’s internal traits. To address this issue, researchers reformat test content or dynamically generate novel tests; see § 6.2 for more details.
Invalid Novel Items.
The reformatting or dynamic generation of novel tests presents new challenges for content validity. The novel test items may capture only limited dimensions of the construct or involve extraneous factors, thus inadequately representing target constructs. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Content validity assessments for custom-curated and model-generated items are crucial but rarely conducted in LLM psychometrics.
@@ 7.2.2 Construct Validity
Construct validity assesses whether a test accurately measures the theoretical construct it is intended to evaluate. For example, a position paper by Alaa et al. [2025] underscores both the deficiency and critical importance of construct validity in medical LLM benchmarks. Broader literature suggests that construct validity could be undermined by the unique internal abstractions, systematic response patterns, and social desirability bias of LLMs.
@@
LLMs and humans may differ fundamentally in how they internally represent psychological constructs. For instance, Kovač et al. [2023] argue that LLMs exhibit contextually adaptive traits rather than the stable psychological characteristics found in humans. Sühr et al. [2023] systematically challenge the validity of personality tests for LLMs. They find personality types do not follow the clear separation into big five traits observed in humans. Similarly, Peereboom et al. [2024] examine the latent personality structures of humans and three LLMs, concluding that the established HEXACO tests for humans are not applicable to LLMs, as the relevant personality factors may be absent in these models. In measuring values, Ye et al. [2025a] find that self-report results are inconsistent with Schwartz’s theoretical model, and that measuring open-ended responses improves the alignment, but not towards a perfect fit. Hadar-Shoval et al. [2024] also agree that LLMs encode values distinct from those of humans. There is a clear consensus among researchers that a robust evaluation and understanding of LLM psychology necessitates new operational definitions that are distinct from, or only loosely analogous to, human traits.
Pioneering research presents value systems specific to LLMs, in response to their unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024]. Notably, Ye et al. [2025b] propose a theoretically grounded generative psycho-lexical approach for system construction based on GPV value measurement [Ye et al., 2025a]. They evaluate their proposed value systems through statistical analysis and explanatory, predictive, and alignment tasks, establishing the validity of the value factors and their superiority over Schwartz’s value system. The unique constructs of LLMs necessitate the rigorous development of tailored tests. Ma et al. [2025] introduce the Core Sentiment Inventory (CSI), an instrument for measuring the emotional tendencies and personalities of LLMs. Fang et al. [2024] curate a psychologically grounded benchmark for mathematics proficiency, aiming for better validity, modeling item difficulty, and allowing for comparison against human norms. Similar attempts are made by Lee et al. [2024d], who propose the TRAIT test, which yields higher validity and reliability than other personality tests.
@@
Response set refers to the systematic pattern of answering questions in a particular way, regardless of the item’s content. Some LLMs demonstrate option position bias towards responses labeled as "A", particularly in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d]. After adjusting for the option position bias, Dominguez-Olmedo et al. [2024] additionally find that LLM survey responses to political opinion questions are uniformly random. In this case, it becomes meaningless to analyze the distributional alignment between LLM responses and opinions of demographic groups. Similar results are observed in LLM personality tests [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a]. Sühr et al. [2023] discover LLMs’ tendency to respond affirmatively to reverse-coded items (e.g., "I am introverted" vs. "I am extroverted"). Ye et al. [2025a] identify that certain LLMs lean towards low/high ratings. They conclude that self-report tests are not the correct tools for LLM psychometrics.
@@
Social desirability bias is the tendency to give responses that conform to social norms rather than reflect true beliefs. Studies indicate that LLMs exhibit this bias similarly to humans. Salecha et al. [2024] report that in personality tests, LLMs tend to skew their responses towards socially favorable traits, leading to higher extraversion scores and lower neuroticism ratings. Likewise, Ye et al. [2025b] find that when LLMs are directly asked to self-report values, as per Biedma et al. [2024], they avoid reporting less socially desirable values like hedonism, although they express these values when given indirect, contextual prompts. While fine-tuning for social desirability is important, it is crucial to maintain a balanced representation of diverse perspectives and values. Over-optimizing for social desirability can result in homogeneous models that fail to capture the full range of human values and do not effectively meet diverse user needs.
@@
Some research investigates the psychometric validity in cross-lingual settings. Romero et al. [2024] administer a standardized personality questionnaire to LLMs in nine languages and find notable inconsistencies in measurement results. Despite their multilingual abilities, LLMs may lack consistency in psychological traits across languages [Cahyawijaya et al., 2024]. These inconsistencies complicate the psychometric validation of cross-lingual tests, necessitating the disentanglement of intrinsic model biases, translation equivalence issues, and cultural differences represented by the models.
@@ 7.2.3 Criterion and Ecological Validity
Criterion validity involves the correspondence of test results with external standards, while ecological validity evaluates the applicability of these outcomes to real-world scenarios. In LLM psychometrics, external standards often overlap with real-world evaluation results. Ren et al. [2024], Ye et al. [2025a] report that value orientations measured by forced-choice tests do not align with those measured in human-LLM interactions. Similar discrepancies are noted in personality [Ai et al., 2024] and morality [Nunes et al., 2024]. Consequently, Zhang et al. [2024a] advocate for linking machine personality assessments to safety. Ren et al. [2024] recommend measuring values within real-world human-LLM interactions, and Ye et al. [2025a] further demonstrate that values measured in open-ended dialog enhance construct validity given Schwartz’s circumplex model. Further work by Ye et al. [2025b] connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards.
@@ 7.3 Standards and Recommendations
Based on the identified challenges and solutions, researchers have proposed general standards and recommendations for LLM psychometrics, aiming to establish a methodological foundation for this nascent field. Frank [2023a] explores how developmental psychology helps us understand the internal representations and cognitive capacities of LLMs. The author recommends using simplified and novel stimuli to avoid the effects of training data contamination. [Löhn et al., 2024] critique the current state of psychometric evaluation on LLMs and argue for standardized criteria to ensure valid assessments. They propose seven requirements: reliability (consistent results), validity (measuring intended traits), suitability (matching LLM capabilities), non-disclosure (avoiding test contamination), and fairness; fairness further includes valid comparisons, valid translations, and transparency. Analyzing 25 studies, they find widespread neglect of these principles.
More recently, Hagendorff et al. [2024] recommend procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options in multiple-choice questions, and using multiple scoring methods to avoid data contamination and ensure reliability. They also suggest using deterministic settings for reproducibility, automated tools for evaluation, and manual reviews for unreliable outputs. After evaluations, they advocate for performing statistical analysis to interpret the results. Vaugrante et al. [2024] identify the lack of replicability in recent works and propose four recommendations. First, they advise ensuring benchmark validity, providing adequate tasks for statistical analysis, standardizing for comparability, controlling prompt sensitivity, and aligning with research objectives. Second, adopt standardized methodologies, avoid cherry-picking, ensure statistical transparency, document experimental setups comprehensively, and define consistent evaluation metrics to enable reliable and replicable research. Third, monitor model behavior changes, account for variability with diverse models, adjust benchmark difficulty as models improve, and ensure transparency by documenting model versions and experiment dates. Lastly, standardize scoring and verification processes, ensuring accuracy and transparency; benchmark creators should provide clear verification guidelines and scoring rubrics for implementers. Schelb et al. [2025] release a framework for designing and running robust and reproducible psychometric experiments. The framework focuses on robustness, flexibility, usability, and reproducibility, based on a standardized configuration file defining experiments.
@@ 8 Psychometrics for LLM Enhancement
Psychometric principles serve as powerful tools for LLM development and enhancement, beyond mere evaluation. This shift has led to three key advancements: trait manipulation for personalized and controlled model behaviors, safety and alignment through psychometrically-informed interventions, and cognitive enhancement to foster more human-like reasoning and communication capabilities.
@@ 8.1 Trait Manipulation
One of the major applications of psychometrics is to manipulate the traits of LLMs, for chatbot personalization, role-play applications, and simulating demographic groups. Psychometrics provides a robust framework to inform such manipulation across prompting, inference, and training stages.
@@
Structured prompting, grounded in validated psychometric inventories, are shown to reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024]. This approach facilitates the controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024]. For example, Jiang et al. [2023] introduce personality prompting ($P^{2}$) based on the lexicon theories behind the Big Five personality traits. In $P^{2}$, lexicons related to specific traits are expanded into detailed personality descriptions by LLMs. The LLMs are then personalized towards diverse behaviors. Chuang et al. [2024] propose a more systematic prompting strategy based on empirically-derived human belief networks, which encompasses 64 topics loading on 9 non-overlapping latent factors. When prompted with one belief, the LLMs become more aligned with other related beliefs.
@@
Beyond prompt engineering, inference-time interventions directly manipulate hidden representations during forward pass, with many based on patterns recognized using psychometric scales. ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] manipulate model outputs on trait axes by directly shifting activations or targeted neuron values at inference time, achieving trait control without retraining.
@@
Other works further fine-tune the LLMs for trait manipulation. Vu et al. [2024] adapt the model architecture to incorporate trait control, reflecting continuous Big Five and mental health dimensions. Dan et al. [2024], Jain et al. [2024], Li et al. [2024c] deploy LoRA modules or routing networks tied to personality traits, learning specialist adapters for trait-driven generation. Cui et al. [2023], Zeng et al. [2024b], Liu et al. [2024a] achieve deep trait embedding by training on large, annotated dialogue datasets derived from validated psychometric instruments. Additional work directly modulates the model parameters without retraining. Hwang et al. [2025] generate adjustment queries based on the MBTI questionnaire and edit the model parameters to align their responses with the desired personality traits.
@@
The manipulation of traits in LLMs presents a promising avenue for substituting human participants in social science research, as reviewed by Gao et al. [2024]. Cao et al. [2025], Suh et al. [2025] demonstrate that fine-tuning LLMs enables them to accurately simulate survey response distributions of specific subpopulations, even when faced with previously unseen survey questions. He-Yueya et al. [2024] introduce the concept of psychometric alignment to evaluate the congruence between LLMs’ knowledge and human understanding. By training LLMs on human response data from the target distribution, they enhance the models’ psychometric alignment on novel test items. However, the efficacy of this training approach is contingent upon the domain. Furthermore, Kang et al. [2023], Sorensen et al. [2025] explore the representation of human values within LLMs and fine-tune them to simulate human opinions through value injection.
@@
Although current personalization techniques have demonstrated the ability to effect more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains a significant challenge [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
@@ 8.2 Safety and Alignment
Recent research has established a connection between LLM psychometric measurement and their safety and alignment, one of the most pressing issues in the field. Zhang et al. [2024a], Wang et al. [2025] link LLM personality traits to their safety. Specifically, Zhang et al. [2024a] identify a significant relationship between personality traits, as measured by the MBTI-M scale, and safety, noting that better alignment enhances traits such as Extraversion, Sensing, and Judging. They are inspired to modify LLM personalities, resulting in improved safety, particularly in privacy and fairness. Similarly, Wang et al. [2025] examine the correlation between HEXACO personality traits and LLM safety.
In parallel, Ye et al. [2025a, b], Yao et al. [2024] have explored the relationship between LLM values and their safety and alignment. Yao et al. [2024] demonstrate the feasibility of distinguishing between safe and unsafe LLM responses within the basic value space, analyzing correlations with Schwartz’s values and specific safety issues. Ye et al. [2025a, b] achieve high accuracy in predicting LLM safety scores based on value orientations measured by GPV, a generative psychometric tool. Furthermore, Yao et al. [2024], Ye et al. [2025b] employ reinforcement learning to align LLMs with desired human values, thereby enhancing their safety. In the domain of morality, Huang et al. [2024a], Tlaie [2024] introduce prompting techniques grounded in Moral Foundation Theory to enhance the moral reasoning and alignment of LLMs. Takeshita et al. [2023], Ohashi et al. [2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to local culture.
@@ 8.3 Cognitive Enhancement
Psychometrics has proven to be an effective tool for developing human-like reasoning, empathy, and communication skills in LLMs. Recent studies leverage established psychological theories and psychometric frameworks to guide LLMs’ cognitive enhancement, employing strategies like prompting, architectural modules, and specialized training regimes.
Psychology-inspired prompting is shown to enhance the cognitive capabilities of LLMs. For example, Li et al. [2023a] introduce emotional prompts, demonstrating that such stimuli can improve the general cognitive abilities of LLMs. Role-playing prompts are observed to influence ToM capabilities [Tan et al., 2024] and promote more human-like reasoning [Nighojkar et al., 2025]. Zhao et al. [2025b] employ a finite-state machine paradigm based on Hill’s Helping Skills theory. It embeds state transitions for emotional support conversations directly within the model’s multi-hop inference process as additional contexts. This method effectively structures dialogues around psychological support strategies and emotional states, leading to higher human-rated effectiveness and strategic alignment.
Liu et al. [2023] introduce a neural listener module within an LLM architecture, optimizing a combined objective that encodes ToM reasoning into the model’s training. They emphasize the potential of using psycholinguistic theories of human language acquisition to enhance LLM language learning. Additionally, preference-based and RL-based methods are proposed to internalize empathetic capabilities in LLMs. Research in empathetic response generation often utilizes cognitive and affective models of empathy to create composite reward functions or preference signals for fine-tuning model behavior. For instance, Sotolar et al. [2024] generate training data using formal emotion theories, like Plutchik’s wheel of emotions, to create preference pairs for model optimization. Others have introduced empathy classifiers to shape reinforcement learning rewards, guiding models toward empathetic outputs [Sharma et al., 2021]. Further research improves LLM social-pragmatic reasoning through preference optimization based on human judgments in open-ended social tasks, showing that these objectives enhance pragmatic inference [Wu et al., 2024].

### v2

@@ 6.2 Validity
Validity assesses whether a test truly measures its intended construct. Recent work in LLM Psychometrics has explored various facets of validity, highlighting both challenges and emerging solutions.
@@ 6.2.1 Content validity
Content validity requires that a test fully represents the intended construct. Key challenges include data contamination from established tests and insufficiently vetted novel items. Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [Hagendorff et al., 2024, Jiang et al., 2024c]. This can inflate performance or bias trait measurement. To mitigate this, researchers reformat tests or generate new items; see § 5.2. On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Evaluating content validity for custom-curated and model-generated items is crucial but rarely conducted in LLM Psychometrics.
@@ 6.2.2 Construct validity
Construct validity determines if a test truly measures its intended construct. Alaa et al. [2025] highlight its critical yet often lacking role in medical LLM benchmarks. Broader literature suggests that construct validity is challenged by LLMs’ unique internal abstractions, systematic response patterns, and social desirability bias.
@@
LLMs and humans may fundamentally differ in their internal representations of psychological constructs. Kovač et al. [2023] argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans. Sühr et al. [2023] show that LLM personality types do not align with the clear separation of Big Five traits found in humans. Similarly, Peereboom et al. [2024] find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent. For values, Ye et al. [2025a] report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; Hadar-Shoval et al. [2024] also find LLMs encode values distinct from humans. Overall, researchers agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
In this regard, pioneering research presents value systems tailored to LLMs’ unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024]. Ye et al. [2025b] introduce a generative psycho-lexical approach based on GPV value measurement [Ye et al., 2025a], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system. The distinct nature of LLM constructs requires rigorously designed tests: Ma et al. [2025] present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; Fang et al. [2024] create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and Lee et al. [2024d] propose the TRAIT test, achieving higher validity and reliability than other personality measures.
@@
Response set refers to systematic answer patterns regardless of item content. Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d]. After correcting for this bias, Dominguez-Olmedo et al. [2024] find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless. Similar findings are observed in LLM personality [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a]. Sühr et al. [2023] also note LLMs often affirm both the original and reverse-coded items, while Ye et al. [2025a] observe some LLMs consistently prefer low or high ratings. These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
@@
Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans. Salecha et al. [2024] find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores. Similarly, Ye et al. [2025b] show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [Biedma et al., 2024], but reveal them in indirect, contextual prompts. While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
@@
Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [Romero et al., 2024]. Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [Cahyawijaya et al., 2024]. These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.
@@ 6.2.3 Criterion and ecological validity
Criterion validity examines how well test results correspond to external standards, while ecological validity concerns their relevance to real-world scenarios. In LLM Psychometrics, external standards often coincide with real-world evaluation results. Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [Ren et al., 2024, Ye et al., 2025a], with similar gaps in personality [Ai et al., 2024] and morality [Nunes et al., 2024]. As a result, researchers advocate linking LLM personality to safety [Zhang et al., 2024a], measuring values within real-world interactions [Ren et al., 2024], and using open-ended dialog to improve construct validity [Ye et al., 2025a]. Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [Ye et al., 2025b].
@@ 6.3 Standards and recommendations
To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor. Frank [2023a] suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination. Löhn et al. [2024] critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency. Their review of 25 studies reveals these principles are often overlooked.
More recently, Hagendorff et al. [2024] advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability. They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results. Vaugrante et al. [2024] highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers. Schelb et al. [2025] introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
@@ 7 Psychometrics for LLM enhancement
Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement. They underpin three main advancements: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
@@ 7.1 Trait manipulation
Psychometrics enables targeted manipulation of LLM traits for personalization, role-play, and demographic simulation, providing a principled framework across prompting, inference, and training.
@@
Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024], supporting controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024]. For instance, Jiang et al. [2023] introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors. Chuang et al. [2024] propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
@@
Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales. Methods such as ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] achieve trait control by shifting activations or neuron values at inference, without retraining.
@@
Other approaches fine-tune LLMs for trait manipulation. Vu et al. [2024] modify model architectures to enable control over continuous Big Five and mental health traits. Dan et al. [2024], Jain et al. [2024], Li et al. [2024c] deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation. Cui et al. [2023], Zeng et al. [2024b], Liu et al. [2024a] embed traits by training on large, annotated dialogue datasets from validated psychometric tools. Some methods directly edit model parameters without retraining; for example, Hwang et al. [2025] use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
@@
Trait manipulation in LLMs offers a promising alternative to human participants in social science research [Gao et al., 2024, Cui et al., 2025]. Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [Cao et al., 2025, Suh et al., 2025]. He-Yueya et al. [2024] propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain. Additionally, Kang et al. [2023], Sorensen et al. [2025] investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
@@
While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
@@ 7.2 Safety and alignment
Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field. Zhang et al. [2024a], Wang et al. [2025b] show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety. For example, Zhang et al. [2024a] find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
Similarly, studies have connected LLM values to safety and alignment. Yao et al. [2024] distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while Ye et al. [2025a, b] accurately predict LLM safety scores from value orientations measured by the GPV tool. Both Yao et al. [2024], Ye et al. [2025b] use reinforcement learning to align LLMs with human values, improving safety. In morality, Huang et al. [2024a], Tlaie [2024] apply MFT-based prompting to enhance LLM moral reasoning and alignment. Takeshita et al. [2023], Ohashi et al. [2024] introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
@@ 7.3 Cognitive enhancement
Psychometrics effectively enhances LLMs’ human-like reasoning, empathy, and communication skills. Recent work applies psychological theories and psychometric frameworks to improve LLM cognition via prompting, architectural modules, and specialized training.
Psychology-inspired prompts boost LLM cognitive abilities. For instance, Li et al. [2023a] show that emotional prompts enhance general cognition. Role-playing prompts improve ToM [Tan et al., 2024] and foster more human-like reasoning [Nighojkar et al., 2025]. Zhao et al. [2025b] propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference. This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
Liu et al. [2023] introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition. Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals. For example, Sotolar et al. [2024] utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [Sharma et al., 2021]. Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [Wu et al., 2024].